# Sub-Riemannian Geometry, Mixing, and the Holonomy of Optimal Mass Transport

### v1

@@ Abstract
The theory of Monge-Kantorovich Optimal Mass Transport (OMT) has in recent years spurred a fast developing phase of research in stochastic control, control of ensemble systems, thermodynamics, data science, and several other fields in engineering and science. Specifically, OMT endowed the space of probability distributions with a rich Riemannian-like geometry, the Wasserstein geometry and the Wasserstein $\mathcal{W}_{2}$-metric. This geometry proved fruitful in quantifying and regulating the uncertainty of deterministic and stochastic systems, and in dealing with problems related to the transport of ensembles in continuous and discrete spaces.
We herein introduce a new type of transportation problems. The salient feature of these problems is that particles/agents in the ensemble, that are to be transported, are labeled and their relative position along their journey is of interest. Of particular importance in our program are closed orbits where particles return to their original place after being transported along closed paths. Thereby, control laws are sought so that the distribution of the ensemble traverses a closed orbit in the Wasserstein manifold without mixing. This feature is in contrast with the classical theory of optimal transport where the primary object of study is the path of probability densities, without any concern about mixing of the flow, which is expected and allowed when traversing curves in the Wasserstein space.
In the theory that we present, we explore a hitherto unstudied sub-Riemannian structure of Monge-Kantorovich transport where the relative position of particles along their journey is modeled by the holonomy of the transportation schedule. From this vantage point, we discuss several other problems of independent interest.
Keywords
Stochastic control, Monge-Kantorovich transport, sub-Riemannian geometry, isoholonomic problem, mixing
@@ 1 Introduction
The history of Optimal Mass Transport (OMT) can be traced to the work of Gaspar Monge in the latter part of the 18th century, who apparently was inspired by practical problems that he encountered in leveling dirt roads [1, 2, 3]. It can be argued that Monge’s problem constitutes one of the first optimal control problems, as he sought a transport schedule, for mass distributed in one location to be moved to another, that accrues minimal cost [4, 5]. Historically, Monge’s cost was the actual Euclidean distance traversed by mass particles, although there are strong reasons to consider a (convex) cost, such as the square of the distance. Furthermore, Monge was interested in path planning besides the end to end correspondence. The history from there on was long, with the name of Kantorovich standing out –he invented duality theory to solve Monge’s problem and received the Nobel prize for the transformative impact of his work.
The modern phase of OMT began in the 1990’s with contributions by McCann, Gangbo, Benamou, Brenier, Otto, and many others [6, 7, 8, 9]. The form of the optimization problem that is pertinent herein, due to Benamou and Brenier [2, Ch. 6], is:
@@
[…]
@@
The starting point of our inquiry is the observation that when the distribution of an ensemble of particles traverses a closed orbit in $\mathcal{P}_{2}(\mathbb{R}^{n})$ before it returns to its starting value, the particles do not necessarily return to their original positions. Indeed, traversing a closed path by an ensemble of particles, as a distribution, typically incurs mixing. For instance, when considering a closed path of three McCann geodesics in $\mathcal{P}_{2}(\mathbb{R}^{n})$ that connect pairs of $\bm{\mu}_{1},\bm{\mu}_{1},\bm{\mu}_{3}$ in succession, unless the Monge maps $\varphi_{i,j}^{\star}$ taking $\bm{\mu}_{j}$ to $\bm{\mu}_{i}$ commute, the composition
@@
The point is further exemplified in Fig. 1 by the rudimentary case of two “tracer” particles (equivalently, Dirac distributions) being transferred in succession between three specified configurations where they are in close proximity to one another, in a way that minimizes the overall Wasserstein distance being traversed. It is observed that after one cycle their positions are exchanged. It is clear that in such cases where distributions consist of finitely many particles (singular distributions), closed orbits induce a permutation of the particles and it is the group of permutations that effects mixing. For more general cases of probability measures in $\mathbb{R}^{n}$ or on a manifold, the natural holonomy groups are measure-preserving diffeomorphisms [10]. The present paper will be devoted to Gaussian probability distributions. We will study optimal transport on the fibre bundle of transportation plans where a holonomy group acts naturally, and we will focus on developing a suitable differential geometric framework to address questions of isoholonomy and mixing.
Specifically, in Section 2 we provide a brief expository on optimal transport of Gaussian distributions highlighing the issue of holonomy that arises when traversing a closed path in the Wasserestein space. In Section 3 we develop the differential geometric setting that is suitable for studying holonomy and parallel transport along Wasserstein curves. In Section 4 we detail the sub-Riemannian structure of the optimal transport, and address the optimal transport problem that realizes a specified parallel transport map and holonomy, the isoparallel mass transport problem. Section 5 overviews examples and applications that highlight the relevance of the theory we presented. We close with an epilogue, in Section 6, where we envision enabling new aspects of the framework and concepts that we discussed in the paper.
@@ 2.2 The Holonomy of Gaussian triangles
$\Sigma_{t}$ :=$:=\begin{cases}(I+3t(\Phi_{1\mapsto 2}^{\star}-I)){\sharp}\mathbf{\Sigma}_{1},&t\in\left[0,\frac{1}{3}\right),\\ (I+3(t-\frac{1}{3})(\Phi_{2\mapsto 3}^{\star}-I)){\sharp}\mathbf{\Sigma}_{2},&t\in\left[\frac{1}{3},\frac{2}{3}\right),\\ (I+3(t-\frac{2}{3})(\Phi_{3\mapsto 1}^{\star}-I)){\sharp}\mathbf{\Sigma}_{3},&t\in\left[\frac{2}{3},1\right],\\ \end{cases}$ (14)
@@
as can be readily verified using (11a). But $\Theta\neq I$, in general. Thus, $X_{1}\neq\Theta X_{0}$ for a tracer particle starting at $X_{0}=x$. The two-dimensional sketch of Figure 2 highlights the point. Following fluid verbiage, we refer to this as “mixing”.
@@
Figure 2: Trajectories of two tracer particle traversing a Gaussian triangle with vertices $\{\mathbf{\Sigma}_{1},\mathbf{\Sigma}_{2},\mathbf{\Sigma}_{3}\}$. The arrows indicate the direction of motion of the tracer particles. Each dashed ellipse represents a level set of the corresponding probability density.
@@ 4.1 The holonomy of linear OMT
Proof: The proposition follows directly from the Chow-Rashevskii theorem, provided that the horizontal sub-bundle is bracket-generating, i.e. satisfies the Hörmander condition [17]. The latter assertion follows from the observation that if $A,B\in\text{Sym}(n)$, then the commutator of the two vector fields $\dot{\Phi}^{A}_{t}:=A\Phi_{t}$ and $\dot{\Phi}^{B}_{t}=B\Phi_{t}$ is
$[\dot{\Phi}^{A}_{t},\dot{\Phi}^{B}_{t}]=(BA-AB)\Phi_{t},$ (36)
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the Lie-bracket. Any matrix $M\in\mathbb{R}^{n\times n}$ can be written as the sum of its symmetric part, say, $A$ and its skew symmetric part $S$. In turn, any skew symmetric part can be written as a finite sum of the commutators of symmetric matrices, i.e.,
$S=\sum_{i=1}^{m}B_{i}C_{i}-C_{i}B_{i},$ (37)
for some $m\in\mathbb{N}$ and $B_{i},C_{i}\in\text{Sym}(n)$. Consequently,
$\text{span}_{\mathbb{R}}\{A,BC-CB\}_{A,B,C\in\text{Sym}(n)}=\mathbb{R}^{n\times n}.$ (38)
Starting from any $\Phi\in\text{GL}^{+}(n)$, it follows that
@@
which implies that the horizontal sub-bundle is, indeed, bracket-generating. $\Box$
As noted, the immediate consequence of Proposition 4 is that $\text{Hol}(\Sigma)=\text{SO}(n,\Sigma_{\text{ref}})$ for every $\Sigma\in\text{Sym}^{+}(n)$. Nevertheless, Proposition 4 goes further by establishing one of the essential ingredients for defining a sub-Riemannian metric on the principal bundle $\pi:\text{GL}^{+}(n)\rightarrow\text{Sym}^{+}(n)$; namely that the horizontal sub-bundle (26) is bracket-generating. The remaining ingredient is a choice of a Riemannian metric on the base space $\text{Sym}^{+}(n)$ of the bundle. Naturally, the most pertinent choice in the context of OMT is the Wasserstein-Otto metric [9, 10].
@@ 4.2 The sub-Riemannian structure of linear OMT
Otto’s metric (4), also known as the Wasserstein-Otto metric [10], specialized to Gaussian distributions takes the form
@@
It is a standard result that a sub-Riemannian metric induces a distance function [17, Definition 3.30]. So let $d_{SR}:\text{GL}^{+}(n)\times\text{GL}^{+}(n)\rightarrow\mathbb{R}_{\geq 0}$ be this sub-Riemannian distance induced by $\mathcal{G}$. A sub-Riemannian geodesic is a horizontal curve
@@
$d_{SR}(\Phi_{0},\Phi_{1})=\int_{0}^{1}{\mathcal{G}_{\Phi_{t}}(\dot{\Phi}_{t},\dot{\Phi}_{t})}^{\frac{1}{2}}\,\text{d}t.$ (43)
There are two types of sub-Riemannian geodesics: normal and abnormal [17, Definition 3.60]. Normal geodesics are differentiable and satisfy the first-order (Pontryagin) necessary conditions for optimality. Abnormal geodesics, on the other hand, do not satisfy the first-order necessary conditions of optimality. As shown by Montgomery [19, Chapter 3, Theorem 3.4], the presence of abnormal geodesics in sub-Riemannian geometry can be a generic phenomenon. If present, abnormal geodesics significantly complicate the local geometry of a sub-Riemannian manifold [17, Chapter 11]. Fortunately, the OMT sub-Riemannian structure is special and does not admit abnormal geodesics. The following proposition establishes this fact. The proof is straightforward and is sketched after the proposition.
Proposition 6
@@
Proof: From (39), we see that only first-order Lie brackets are required to generate the tangent bundle from the horizontal sub-bundle of the OMT connection. This property appears under many names in the literature. In [17], the horizontal sub-bundle is said to be a “step-2” bundle when this property holds. Elsewhere, the property has been referred to as the strong bracket generating condition [25]. Strichartz established in [25, Theorem 6.1, Corollary 6.2] that, under the strong bracket generating condition, all sub-Riemannian geodesics are normal, see also [17, Corollary 12.14]. Thus, the same is true for the OMT sub-Riemannian structure. $\Box$
@@
Proof: The Hamiltonian associated with the geodesics of the OMT sub-Riemannian structure is given in coordinates by
@@
The proof is concluded by noting that ${\Phi}_{t}\Sigma_{\text{ref}}{\Phi}_{t}^{\intercal}=\pi(\Phi_{t})$, and that the solution to (46) is expressed via the linear operator defined in (31). $\Box$
@@
Proof: A principal bundle isomorphism [21, p. 53] between $\pi$ and $\tilde{\pi}$ is a diffeomorphism $\Psi:\text{GL}^{+}(n)\rightarrow\text{GL}^{+}(n)$ and a group isomorphism $\psi:\text{SO}(n,\Sigma_{\text{ref}})\rightarrow\text{SO}(n,\tilde{\Sigma}_{\text{ref}})$ such that
@@
It follows that $\Psi$ is also an isometry [25, Theorem 8.2]. $\Box$
@@
Inspection of (42) reveals that this is precisely the restriction to the horizontal sub-bundle of a $\Sigma_{\text{ref}}$-weighted Frobenius inner product on each tangent space $T_{\Phi}\text{GL}^{+}(n)$. Extending (42) to the entire tangent bundle defines a (flat) Riemannian metric, henceforth referred to as the $L_{2}$-metric, on $\text{GL}^{+}(n)$. In fact, the Wasserstein-Otto metric (40) is typically constructed by showing that the map $\pi$ is a Riemannian submersion and that the $L_{2}$-metric descends to the Wasserstein-Otto metric (40) under $\pi$ [9, 24]. This observation is crucial in establishing many important regularity properties of the sub-Riemannian distance $d_{SR}$. For instance, the distance $d_{SR}$ induces the same topology as its Riemannian extension. Another consequence is that, under the strong bracket generating condition, the distance $d_{SR}$ is bounded above on compact sets by a multiple of the square root of its Riemannian extension [25, Theorem 11.1]. This important fact is recapitulated in the following proposition.
Proposition 9
@@
Figure 3: Trajectories of three tracer particle traversing the McCann geodesic connecting $\Sigma_{\text{in}}$ and $\Sigma_{\text{fn}}$ prescribed in (58). The dashed black ellipses mark intermediate covariances along the trajectory.
Figure 4: Trajectories of three tracer particle traversing the isoparallel curve connecting $\Sigma_{\text{in}}$ and $\Sigma_{\text{fn}}$ prescribed in (58) with $\Theta$ defined in (59). The dashed black ellipses mark intermediate covariances along the trajectory.

### v2

# The Holonomy of Optimal Mass Transport: The Gaussian-Linear Case

@@ Abstract
The theory of Monge-Kantorovich Optimal Mass Transport (OMT) has in recent years spurred a fast developing phase of research in stochastic control, control of ensemble systems, thermodynamics, data science, and several other fields in engineering and science.
We herein introduce a new type of transportation problems. The salient feature of these problems is that particles/agents in the ensemble are labeled and their relative position along their journey is of interest. Of particular importance in our program are control laws that steer ensembles along cycles ensuring that individual particles return to their original position. This feature is in contrast with the classical theory of optimal transport where the primary object of study is the path of probability densities, without any concern about particle labels.
In the theory that we present, we focus on the case Gaussian distributions and linear dynamics, and explore a hitherto unstudied sub-Riemannian structure of Monge-Kantorovich transport where the relative position of particles along their journey is modeled by the holonomy of the transportation schedule. From this vantage point, we discuss several other problems of independent interest.
Keywords Monge-Kantorovich transport, sub-Riemannian geometry, mixing
@@ 1 Introduction
The history of Optimal Mass Transport (OMT) can be traced to the work of Gaspar Monge in the latter part of the 18th century, who apparently was inspired by practical problems that he encountered in leveling dirt roads [37, 43, 44, 11]. It can be argued that Monge’s problem constitutes one of the first optimal control problems, as he sought a transport schedule, for mass distributed in one location to be moved to another, that accrues minimal cost [10]. Historically, Monge’s cost was the actual Euclidean distance traversed by mass particles, although there are strong reasons to consider a (convex) cost, such as the square of the distance. Furthermore, Monge was interested in path planning besides the end to end correspondence. The history from there on was long, with the name of Kantorovich standing out –he invented duality theory to solve Monge’s problem and received the Nobel prize for the transformative impact of his work.
@@
and terminal conditions at $t\in\{0,1\}$; $n$ is the dimension of the space and $\nabla\cdot$ denotes the divergence. Here, $u(t,\cdot)$ represents a vector field (our control action) that steers an initial probability density [^1] Following a common slight abuse of notation one often uses the same symbol $\mu$ to denote the corresponding measure $\mu(dx)$, that in our case is always assumed absolutely continuous with respect to the Lebesgue. $\mu_{0}(x)$ to a final one $\mu_{1}(x)$ over $t\in[0,1]$.
@@
The starting point of our inquiry is the observation that when the distribution of an ensemble of particles traverses a closed orbit in $\mathcal{P}_{2}(\mathbb{R}^{n})$ before it returns to its starting value, the particles do not necessarily return to their original positions. Indeed, traversing a closed path by an ensemble of particles, as a distribution, generically incurs a rearrangement of particle positions. For instance, when considering a closed path of three McCann geodesics in $\mathcal{P}_{2}(\mathbb{R}^{n})$ that connect pairs of $\bm{\mu}_{1},\bm{\mu}_{1},\bm{\mu}_{3}$ in succession, unless the Monge maps $\varphi_{i,j}^{\star}$ taking $\bm{\mu}_{j}$ to $\bm{\mu}_{i}$ commute, the composition
@@
The point is further exemplified in Fig. 1 by the rudimentary case of two “tracer” particles (equivalently, Dirac distributions) being transferred in succession between three specified configurations where they are in close proximity to one another, in a way that minimizes the overall Wasserstein distance being traversed. It is observed that after one cycle their positions are exchanged. Clearly, in such cases where distributions consist of finitely many particles (singular distributions), closed orbits induce a permutation of the particles and it is the group of permutations that effects rearrangement of particle positions. For more general probability measures on $\mathbb{R}^{n}$ or on a manifold, and under sufficient regularity conditions, the relevant groups are measure-preserving diffeomorphisms [21].
Following fluid verbiage, we refer to the phenomenon outlined above, i.e., the rearrangement of particle positions without altering their distribution, as “mixing”. The adopted nomenclature is not to be confused with the classical notion of mixing in ergodic theory where one studies the limiting behavior of repeated compositions of general measure preserving maps. As will be shown in the sequel, the notion of mixing adopted herein turns out to have a natural interpretation as the holonomy [24] of loops in the Wasserstein manifold $\mathcal{P}_{2}(\mathbb{R}^{n})$. The present paper is devoted to the study of said holonomy in the case of Gaussian probability distributions. That is, we study optimal transport in the fibre bundle of linear transportation plans, where a holonomy group acts naturally, and focus on developing a suitable differential geometric framework to study such rearrangement of particle positions.
Specifically, in Section 2 we provide a brief expository on optimal transport of Gaussian distributions highlighting the holonomy that arises when traversing a closed path in the Wasserstein space. In Section 3 we develop the differential geometric setting that is suitable for studying parallel transport and holonomy along Wasserstein curves. In Section 4 we detail a hitherto unstudied sub-Riemannian structure induced by optimal transport wherein the problem of realizing a specified parallel transport map along optimal curves in the Wasserstein space can be naturally formulated and studied. Section 5 overviews examples and applications that highlight the relevance of the theory we presented. We close with an epilogue, in Section 6, where we envision enabling new aspects of the framework and concepts discussed in the paper.
@@ 2.2 The Holonomy of Gaussian triangles
[…]
@@
as can be readily verified using (11a). But $\Theta\neq I$, in general. Thus, $X_{1}\neq\Theta X_{0}$ for a tracer particle starting at $X_{0}=x$. The two-dimensional sketch of Figure 2 highlights the point.
@@
Figure 2: Trajectories of two tracer particle traversing a Gaussian triangle with vertices $\{\mathbf{\Sigma}_{1},\mathbf{\Sigma}_{2},\mathbf{\Sigma}_{3}\}$. The arrows indicate the direction of motion of the tracer particles. Each colored ellipse represents a level set of the corresponding Gaussian probability density.
@@ 4.1 The holonomy of linear OMT
Proof.
The proposition follows directly from the Chow-Rashevskii theorem, provided that the horizontal sub-bundle is bracket-generating, i.e., satisfies the Hörmander condition [1]. The latter assertion follows from the observation that if $A,B\in\text{Sym}(n)$, then the commutator of the two vector fields $\dot{\Phi}^{A}_{t}:=A\Phi_{t}$ and $\dot{\Phi}^{B}_{t}=B\Phi_{t}$ is
$[\dot{\Phi}^{A}_{t},\dot{\Phi}^{B}_{t}]=\bm{[}\hskip-2.0pt\bm{[}B,A\bm{]}\hskip-2.0pt\bm{]}\Phi_{t}=(BA-AB)\Phi_{t},$ (36)
[…]
$A=\bm{[}\hskip-2.0pt\bm{[}B,C\bm{]}\hskip-2.0pt\bm{]},$ (37)
for some $B,C\in\text{Sym}(n)$.
[…]
blackBefore we continue with the proof of the proposition, we give a short proof of the last statement. The validity of the statement can easily be seen in the case of $2\times 2$ skew-symmetric matrix, since any such matrix is of the form
$A_{\lambda}:=\begin{pmatrix}0&\lambda\\ -\lambda&0\end{pmatrix}$,
for some $\lambda\in\mathbb{R}$, while observing that $A_{\lambda}=\bm{[}\hskip-2.0pt\bm{[}\lambda E_{1},E_{2}\bm{]}\hskip-2.0pt\bm{]}$ for
$E_{1}$ $=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$, $E_{2}$ $=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$.
In general, any skew-symmetric matrix $A$ can be brought by an orthogonal transformation $\Theta$ into a block diagonal form
$\Theta A\Theta^{\top}=\begin{bmatrix}A_{\lambda_{1}}&0&\cdots&0\\ 0&A_{\lambda_{2}}&&0\\ \vdots&&\ddots&\vdots\\ 0&0&\cdots&A_{\lambda_{r}}\end{bmatrix}$,
for some $\lambda_{1},\ldots,\lambda_{r}\in\mathbb{R}$ when $n$ is even, or with an additional 0 on the diagonal when $n$ is odd. (This last statement can be easily shown using the fact that any skew-symmetric matrix is normal.) Thus, in general, $A$ can be written as the commutator of the symmetric matrices
$B=\Theta^{\top}\begin{bmatrix}\lambda_{1}E_{1}&0&\cdots&0\\ 0&\lambda_{2}E_{1}&&0\\ \vdots&&\ddots&\vdots\\ 0&0&\cdots&\lambda_{r}E_{1}\end{bmatrix}\Theta,$
and
$C=\Theta^{\top}\begin{bmatrix}E_{2}&0&\cdots&0\\ 0&E_{2}&&0\\ \vdots&&\ddots&\vdots\\ 0&0&\cdots&E_{2}\\ \end{bmatrix}\Theta,$
when $n$ is even, and similarly constructed matrices with an additional 0 on the diagonal if $n$ is odd.
[…]
blackReturning to the proof, we conclude that
$\text{span}_{\mathbb{R}}\{A,\bm{[}\hskip-2.0pt\bm{[}B,C\bm{]}\hskip-2.0pt\bm{]}\}_{A,B,C\in\text{Sym}(n)}=\mathbb{R}^{n\times n}.$ (38)
from whence it follows that, for any $\Phi\in\text{GL}^{+}(n)$,
@@
and, thus, the horizontal sub-bundle is bracket-generating. ∎
[…]
black As an immediate consequence of Proposition 4, we have the following Corollary.
Corollary 1.
$\text{Hol}(\Sigma)=\text{SO}(n,\Sigma_{\text{ref}})$ for all $\Sigma\in\text{Sym}^{+}(n)$.
Proof.
Since Proposition 4 establishes that any two points in $\text{GL}^{+}(n)$ can be connected with a horizontal curve, the same applies to any two points that belong to the same fiber. These are then necessarily related by an element of the structure group, i.e. $\mathrm{SO}(n,\Sigma_{\mathrm{ref}})$. Thus, every $\Theta\in\mathrm{SO}(n,\Sigma_{\mathrm{ref}})$ can be realized as the parallel transport map along a loop in $\text{Sym}^{+}(n)$. The conclusion of the corollary follows. ∎
[…]
blackProposition 4 goes beyond the statement of the Corollary since it establishes one of the essential ingredients for defining a sub-Riemannian metric on the principal bundle $\pi:\text{GL}^{+}(n)\rightarrow\text{Sym}^{+}(n)$; namely that the horizontal sub-bundle (26) is bracket-generating. The remaining ingredient is a choice of a Riemannian metric on the base space $\text{Sym}^{+}(n)$ of the bundle. Naturally, the most pertinent choice in the context of OMT is the Wasserstein-Otto metric [35, 21].
@@ 4.2 The sub-Riemannian structure of linear OMT
Otto’s metric (4), also known as the Wasserstein-Otto metric [21], specialized to Gaussian distributions, takes the form
@@
It is a standard result that a sub-Riemannian metric induces a distance function [1, Definition 3.30]. So let $d_{SR}:\text{GL}^{+}(n)\times\text{GL}^{+}(n)\rightarrow\mathbb{R}_{\geq 0}$ be this sub-Riemannian distance induced by $\mathcal{G}$. A sub-Riemannian geodesic is a horizontal curve
@@
$d_{SR}(\Phi_{0},\Phi_{1})^{2}=\int_{0}^{1}{\mathcal{G}_{\Phi_{t}}(\dot{\Phi}_{t},\dot{\Phi}_{t})}\,\text{d}t.$ (43)
[…]
blackThe regularity of geodesics in sub-Riemannian geometry can be quite subtle. Geodesics are typically characterized as normal and abnormal [1, Definition 3.60]. Normal geodesics are differentiable and satisfy the first-order (Pontryagin) necessary conditions for optimality. On the hand, abnormal geodesics may or may not satisfy the first-order necessary conditions of optimality. Accordingly, there are abnormal geodesics that are also normal, and others that are not; the latter are referred to as strictly abnormal. As shown by Montgomery [32, Chapter 3, Theorem 3.4], the presence of strictly abnormal geodesics can be a generic phenomenon. If present, strictly abnormal geodesics significantly complicate the local geometry of a sub-Riemannian manifold [1, Chapter 11]. Fortunately, the OMT sub-Riemannian structure is special and does not admit strictly abnormal geodesics. The following proposition establishes this fact. The proof is straightforward and is sketched after the proposition.
Proposition 6.
@@
Proof.
From (39), we see that only first-order Lie brackets are required to generate the tangent bundle from the horizontal sub-bundle of the OMT connection. This property appears under many names in the literature. In [1], the horizontal sub-bundle is said to be a “step-2” bundle when this property holds. Elsewhere, the property has been referred to as the strong bracket generating condition [40]. Strichartz established in [40, Theorem 6.1, Corollary 6.2] that, under the strong bracket generating condition, all sub-Riemannian geodesics are normal, see also [1, Corollary 12.14]. Thus, the same is true for the OMT sub-Riemannian structure. ∎
@@
Proof.
The Hamiltonian associated with the geodesics of the OMT sub-Riemannian structure is given in coordinates by
@@
The proof is concluded by noting that ${\Phi}_{t}\Sigma_{\text{ref}}{\Phi}_{t}^{\intercal}=\pi(\Phi_{t})$, and that the solution to (46) is expressed via the linear operator defined in (31). ∎
@@
Proof.
A principal bundle isomorphism [24, p. 53] between $\pi$ and $\tilde{\pi}$ is a diffeomorphism $\Psi:\text{GL}^{+}(n)\rightarrow\text{GL}^{+}(n)$ and a group isomorphism $\psi:\text{SO}(n,\Sigma_{\text{ref}})\rightarrow\text{SO}(n,\tilde{\Sigma}_{\text{ref}})$ such that
@@
It follows that $\Psi$ is also an isometry [40, Theorem 8.2]. ∎
@@
Inspection of (42) reveals that this is precisely the restriction to the horizontal sub-bundle of a $\Sigma_{\text{ref}}$-weighted Frobenius inner product on each tangent space $T_{\Phi}\text{GL}^{+}(n)$. Extending (42) to the entire tangent bundle defines a (flat) Riemannian metric, henceforth referred to as the $L_{2}$-metric, on $\text{GL}^{+}(n)$. In fact, the Wasserstein-Otto metric (40) is typically constructed by showing that the map $\pi$ is a Riemannian submersion and that the $L_{2}$-metric descends to the Wasserstein-Otto metric (40) under $\pi$ [35, 30]. This observation is crucial in establishing many important regularity properties of the sub-Riemannian distance $d_{SR}$. For instance, the distance $d_{SR}$ induces the same topology as its Riemannian extension. Another consequence is that, under the strong bracket generating condition, the distance $d_{SR}$ is bounded above on compact sets by a multiple of the square root of its Riemannian extension [40, Theorem 11.1]. This important fact is recapitulated in the following proposition.
Proposition 9.
@@
Figure 3: Trajectories of three tracer particle traversing the McCann geodesic connecting $\Sigma_{\text{in}}$ and $\Sigma_{\text{fn}}$ prescribed in (58).
Figure 4: Trajectories of three tracer particle traversing the isoparallel curve connecting $\Sigma_{\text{in}}$ and $\Sigma_{\text{fn}}$ prescribed in (58) with $\Theta$ in (59).
Figure 5: Trajectories of three tracer particle traversing the isoholonomic curve connecting $\Sigma_{\text{in}}=I$, $\Sigma_{\text{fn}}=I$, with $\Theta$ in (59).